Given two commodities A and B, where price of A is ₦50 and the price of B is ₦60. If the consumer’s income is ₦1200 and the utility function U = AB, Find: i. the equilibrium values of the two commodities. ii. Mention four properties of the demand function derived in (i) above


1
Expert's answer
2021-02-24T15:02:55-0500
"p_A=50"

"p_B=60"

"I=1200"

"U=AB"

"MU_A=B"

"MU_B=A"

"50A+60B=1200"

"\frac{B}{50}=\frac{A}{60}"

"B=\frac{5}{6}A"

"A=12"

"B=10"